### v1

# Leveraging Generative AI for Smart City Digital Twins: A Survey on the Autonomous Generation of Data, Scenarios, 3D City Models, and Urban Designs

Abstract
The digital transformation of modern cities by integrating advanced information, communication, and computing technologies has marked the epoch of data-driven smart city applications for efficient and sustainable urban management. Despite their effectiveness, these applications often rely on massive amounts of high-dimensional and multi-domain data for monitoring and characterizing different urban sub-systems, presenting challenges in application areas that are limited by data quality and availability, as well as costly efforts for generating urban scenarios and design alternatives. As an emerging research area in deep learning, Generative Artificial Intelligence (AI) models have demonstrated their unique values in data and code generation. This survey paper aims to explore the innovative integration of generative AI techniques and urban digital twins to address challenges in the realm of smart cities in various urban sectors, such as transportation and mobility management, energy system operations, building and infrastructure management, and urban design. The survey starts with the introduction of popular generative AI models with their application areas, followed by a structured review of the existing urban science applications that leverage the autonomous capability of the generative AI techniques to facilitate (a) data augmentation for promoting urban monitoring and predictive analytics, (b) synthetic data and scenario generation, (c) automated 3D city modeling, and (d) generative urban design and optimization. Based on the review, this survey discusses potential opportunities and technical strategies that integrate generative AI models into the next-generation urban digital twins for more reliable, scalable, and automated management of smart cities.
keywords:
Generative artificial intelligence , smart city , urban digital twin , deep learning
1 Introduction
The digital transformation of cities, propelled by advancements in the Internet of Things (IoT), Information and Communication Technologies (ICT), and 5G cellular networks, is leading to the generation of substantial amounts of urban data (Allam and Dhunny, 2019; Javed et al., 2022), enabling the monitoring and situational awareness of various urban sub-systems (Hossain et al., 2018). Driven by the big urban data, smart city applications have reshaped the management of the environment and urbanization in different urban sectors to proliferate the sustainability, productivity, livability, and economic growth in cities across the world (Ullah et al., 2020; Kim et al., 2020a). Examples of these urban sectors include intelligent transportation (Xiong et al., 2012; Zhao and Jia, 2021), urban mobility (Xu et al., 2023b; Paiva et al., 2021), energy smart buildings (Khajenasiri et al., 2017; Francisco et al., 2020), water management and distribution (Karwot et al., 2016), and power grid operations (Omitaomu and Niu, 2021). With the recent advent of Artificial Intelligence (AI) and Digital Twin (DT) technologies, traditional machine learning and deep learning techniques have been ubiquitously applied in smart city applications to support advanced urban monitoring and situational awareness, facilitate complex decision-making, as well as the automation and autonomy in complex urban management processes. (Allam and Dhunny, 2019; Luckey et al., 2021; Cugurullo, 2020; Farsi et al., 2020). Many smart city applications adopt the paradigm of emerging digital twins to develop urban digital twins (i.e., smart city digital twins or digital twin cities) (Shahat et al., 2021; Dembski et al., 2020). The original concept of “Digital Twin” focuses on the development of virtual models of real-world objects and processes in a computing environment to support advanced manufacture and analytics (Ivanov et al., 2020), and was applied to the urban science sector to create urban digital twins, which are digital replicas of cities and their sub-urban system for intelligent urban management and planning, as well as for enhancing smart city servers. These urban digital twins leverage advanced computing and cyberinfrastructure technologies to promote and automate smart city services (Shirowzhan et al., 2020; Xu et al., 2023b). Urban digital twins represent a groundbreaking leap in urban science and planning, especially in the realms of mobility management, building energy optimization, and urban design. By creating detailed, dynamic digital replicas of urban environments, these tools enable planners and scientists to simulate and analyze complex systems with unprecedented accuracy and efficiency in many urban sectors, such as transportation and mobility management (Xu et al., 2023b), building energy systems (Bortolini et al., 2022), smart grids (Cioara et al., 2021), and infrastructure management (Callcut et al., 2021; Sabri et al., 2023). In the broader context of urban design, digital twins offer a holistic view of the urban ecosystem, enabling planners to create more livable, resilient, and adaptive urban spaces (Dembski et al., 2020). The integration of urban digital twins in these sectors not only streamlines planning processes but also fosters examining scenarios and innovative solutions to contemporary urban challenges, paving the way for smarter, more sustainable cities. Some examples are using urban digital twins for water-sensitive urban design (Langenheim et al., 2022), and development envelop control analysis (Sabri et al., 2022).
1.1 Challenges in Urban Digital Twins
Although there have been significant advancements in AI, communication, and computing technologies, urban (or smart city) digital twins that rely on big urban data still encounter various methodological and technical challenges in many aspects:
Data Availability and Quality:
Among the massive smart city data, only a small fraction is utilized by smart services to promote urban management (Mohammadi and Al-Fuqaha, 2018), as much raw data collected from IoT and smart devices, and crowed-sourcing platforms are considered as digital wastes. This is because most of the raw urban data are either unlabeled or have limited data quality, rendering these data unusable for many urban science applications involving data and physical modeling (Park and Ko, 2021; Alwan et al., 2022). Furthermore, a variety of urban data, such as building occupancy, energy consumption, human mobility, and vehicle trajectories, frequently faces limited availability issues owing to data privacy and residence concerns, heterogeneous sources, and nonuniform distribution of sensors (Figueira and Vaz, 2022; Zheng et al., 2014). Subsequently, these datasets exhibit varying coverage, accuracy, and resolution levels across different spatial regions, rendering them less applicable to many analytical applications.
Hypothetical Urban Scenarios:
Many smart city applications involve strategic approaches that envision and analyze various potential future scenarios to enhance urban planning and management (Perveen et al., 2017; Sokolov et al., 2019) in different sectors, such as traffic (Feng et al., 2023a; Xinxin et al., 2020), building energy (Ruan et al., 2023; Pippia et al., 2021), and water management (Herman et al., 2016; Mikovits et al., 2018). These approaches often require costly, time-consuming, and labor-intensive efforts for collecting scenario-specific data, as well as for generating the corresponding scenarios using processed data and simulations to support predictive analytics and decision-making (Dong et al., 2021; Argota Sánchez-Vaquerizo, 2022). In addition, it is also extremely difficult to collect real-world data to describe a future or hypothetical scenario, which are often considered as corner cases.
3D City Modeling:
There is an increasing trend for smart city initiatives to develop 3D city models as critical components of urban digital twins to support data integration and visualization, interactive planning and simulation, real-time monitoring and response, and public engagement (Deng et al., 2021; Farsi et al., 2020). However, developing a comprehensive 3D city model to digitize a large urban area in a computing system can be a costly, repetitive, and labor-intensive endeavor (Botín-Sanabria et al., 2022; Singh et al., 2021), which involves massive modeling and gamification of urban entities (e.g., buildings and infrastructure) as 3D models with different levels of details. Additionally, constructing 3D terrain, underground urban objects, and vegetation using geomatics and surveying data can also be a time-consuming effort (Xu et al., 2023a).
Urban Design:
Built upon the previous data analytics, simulations, and digital twin applications, the planning and management of smart cities often involve top-down and multi-scale urban design processes, which are responsible for the configuration of land-use and land-cover, transportation network, buildings, and infrastructure (Van Nes and Yamu, 2021; Pérez-Martínez et al., 2023). These processes are often repetitive, expensive, and time-consuming, involving multiple design alternatives and complex decision-making for design optimization. Therefore, it is challenging to scale up the design processes to large urban areas, which often have distinct design requirements (e.g., green space and accessibility).
Given the complexities, crafting wide-ranging and insightful solutions to urban issues from an all-encompassing viewpoint is proving to be progressively more expensive and demanding in terms of effort. These solutions require the analysis of complex urban systems — including transportation, energy, building, infrastructure, and water — as interconnected subsystems at the ground, above-ground, and underground levels. The selection of these urban subsystems is based on previous literature and urban digital applications (Ferré-Bigorra et al., 2022, 2022; Hämäläinen, 2021). Consequently, there is a growing need for developing scalable and autonomous solutions. Such solutions would employ generative AI techniques to enhance the availability and quality of urban data, automate the generation of various scenarios, support the development of 3D city models, and automate complex urban design and optimization processes (Wang et al., 2023a; Kirwan and Zhiyong, 2020).
Figure 1: A summary of generative AI-related publications in major urban management sector using Scopus database (Search conducted on Jan 10th 2024)
1.2 Motivation
In the realm of deep learning, Generative Artificial Intelligence (AI) has emerged as a groundbreaking field, showcasing its unique ability to create novel content from realistic data. This capability is increasingly relevant for enhancing smart city applications. Despite the rapid proliferation of generative AI in various urban science and management sectors, there remains a notable absence of a thorough review encompassing mainstreaming generative AI techniques and their contributions to urban and environmental management within the context of developing smart cities through digital twin models. This survey paper aims to fill this gap by providing a structured overview of current generative AI models and their applications in diverse urban management sectors, particularly those not yet explored from the smart city digital twin perspective. Drawing upon recent literature and proposed agendas, the focus of this survey is primarily on key smart city application areas, such as transportation and mobility management, energy system operations, urban water management, and building and infrastructure management (Shahat et al., 2021; Dembski et al., 2020). This survey searched and discovered the publication history and trend of generative AI-powered studies in these application areas using the Scopus database, and present our discovery in Figure 1. In addition, this paper reviews the capability of generative AI techniques in urban analysis (e.g., population and morphology), 3D city modeling, and generative urban design. Organized around the limitations of urban digital twins in data availability and quality, scenario generation, 3D city modeling, and urban design, this survey critically examines the potential integration of generative AI models to enhance the capability, reliability, and efficiency of urban digital twins. Finally, the paper discusses future challenges and opportunities from a software engineering perspective, offering insights into the evolving landscape of smart city development.
2 Generative AI Models
Generative AI models are a class of artificial intelligence techniques designed to generate new data samples similar to a given set of input data (Bandi et al., 2023). These models can produce a wide range of outputs, including images, text, sound, and video, and are particularly notable for their ability to create realistic, novel, and often indistinguishable data from actual human-generated content (Eigenschink et al., 2021), allowing them to support data-driven research application in various urban science and smart city sectors. In addition, Generative AI techniques and services enabled through Large Language Models (LLMs), such as ChatGPT, have demonstrated their unique capabilities in interpreting human natural language and automating code generation and information, therefore, have been applied to develop autonomous systems in various research fields, such as geoinformatics (Li and Ning, 2023), chemical research (Boiko et al., 2023), and transportation (Cui et al., 2024). We discovered the previously published research articles through the Scopus database. We summarized popular generative AI models used in major urban digital twin application areas (as identified in the motivation section). Based on our search combined with summaries proposed by other review articles (Lin et al., 2023; Bandi et al., 2023), this survey paper primarily targets four generative AI models, including the Generative Adversarial Networks, Variational Autoencoders, ransformer-Based Models, and Generative Diffusion Models. These models and their urban applications are briefly and individually discussed in the following subsections.
2.1 Generative Adversarial Networks
Generative Adversarial Networks (GANs) represent a pivotal innovation in the field of artificial intelligence, particularly within the domain of deep learning (Pan et al., 2019b; Goodfellow et al., 2014). These models introduce a novel framework for generative modeling, leveraging the power of two neural networks—a generator and a discriminator—in a competitive setting. The generator aims to produce data indistinguishable from real data, while the discriminator endeavors to distinguish between real and generated data (Goodfellow et al., 2020; Creswell et al., 2018). This adversarial process results in the generator creating increasingly realistic outputs. Over time, GANs have evolved into various forms, such as Conditional GANs, which generate data based on specific conditions (Mirza and Osindero, 2014), and CycleGANs, used for style transfer tasks (Zhu et al., 2017). Their vast and transformative applications range from image and video generation to drug discovery and advanced image editing. GANs have revolutionized how machines understand and recreate complex data patterns, making them a cornerstone technology in modern AI research and development. Their continued evolution promises further breakthroughs in AI’s capability to generate realistic, nuanced, and context-aware content.
Major application areas of GANs in various urban science and smart city sectors include the following:
  • 1.
    Image Generation: GANs can generate highly realistic images, often indistinguishable from actual photographs. This includes creating artwork, enhancing image resolution, or generating images from descriptions. In smart city applications, GANs’ image generation capability is often leveraged to create synthetic urban and environmental datasets in the geospatial raster format. Examples of these synthetic data include river (Gautam et al., 2022), land use (Ansith and Bini, 2021), and urban morphology patterns (Zhang et al., 2022).
  • 2.
    Data Augmentation: the process is used to increase the size and diversity of an existing dataset. GANs can create additional data with improved spatiotemporal resolution for training predictive models, enhancing the robustness of these models without the need to collect more real-world data. In recent years, GANs have been applied to augment urban data to improve the prediction accuracy of many urban variables in areas that include transportation management (Dabboussi, 2023; Lin et al., 2023), building energy optimization (Fan et al., 2022), and infrastructure management (Gwon et al., 2023).
  • 3.
    Synthetic Data for Scenario Generation: the primary goal of synthetic data generation is to create entirely new data points that are not present in the original dataset. This is especially useful when the amount of real data available is limited (e.g., hypothetical scenarios or corner cases) or when using real data is restricted due to privacy concerns (e.g., personal or sensitive information). GANs are employed to generate synthetic data to define many hypothetical scenarios (Sankar, 2020) with a wide range of urban management interests, such as promoting transportation planning (Wu et al., 2020; Demetriou et al., 2020; Rempe et al., 2022), urban mobility management (Wang et al., 2020), power system operations (Yuan et al., 2021; Huang et al., 2023; Rizzato et al., 2022), building energy modeling (Wang and Hong, 2020; Zhang and Zhang, 2019), and water management (Koochali et al., ).
  • 4.
    3D Object Generation: GANs can also generate 3D models that are useful for multiple-purpose visualization functions (e.g., virtual reality, gaming, and filmmaking) and urban science applications. These applications include creating 3D city models (Veselỳ, 2022), 3D buildings (Du et al., 2020a; Kim et al., 2020b), and architectural design (Chang et al., 2021). The GANs’ capability of generation of 3D urban objects has been employed to automate the development of 3D representation of cities in large-scale urban digital twins, which can enable virtual and augmented reality applications for scientific gamification and immersive 3D scientific visualization of urban management processes, shedding light on public engagement and participatory planning in the smart city (Xu et al., 2023a).
2.1.1 Conditional GAN
Conditional Generative Adversarial Networks (cGANs), an influential extension of the original Generative Adversarial Network (GAN) framework, were introduced to add a directed feature to the generative process (Mirza and Osindero, 2014). cGANs were conceptualized to address the limitation of traditional GANs in controlling the modes of the data generation process. Mirza and Osindero (2014), in their seminal work, proposed this adaptation where both the generator and the discriminator are conditioned on some extra information, such as class labels or data from other modalities, allowing the cGANs to generate more specific and targeted outputs with improved relevance and applicability of the generated data (Mirza and Osindero, 2014). The practical use cases of cGANs are vast and diverse, particularly shining in the smart city fields where tailored data generation is crucial. Examples of these fields include urban traffic estimation and predictions (Zhang et al., 2020; Huang et al., 2020), air quality management (Toutouh, 2021), and building operations (Baasch et al., 2021; Năstăsescu and Cercel, 2022). The ability of cGANs to generate contextually relevant and specific data opens new possibilities in various fields, making them a focal point in ongoing AI research.
2.1.2 CycleGANs
CycleGAN, short for Cycle-Consistent Generative Adversarial Networks, represents a significant advancement in the field of machine learning and computer vision. This innovative framework facilitates image-to-image translation tasks without requiring paired training data (Zhu et al., 2017). This is a notable departure from traditional methods that rely heavily on paired datasets, which are often difficult and expensive to obtain. CycleGAN leverages a novel approach that uses two sets of Generative Adversarial Networks (GANs) and a cycle-consistency loss function CycleGAN. This ensures that the inputs closely resemble the original inputs when translated to another domain and then back. This breakthrough has wide-ranging applications, from photo-realistic image translation to style transfer and even in tasks such as object transfiguration and season transfer in images. Its ability to learn mappings between different domains without paired examples makes CycleGAN a powerful tool in unsupervised learning, offering new possibilities in various smart city fields such as autonomous driving (Rajagopal et al., 2023), augmented reality and virtual reality (Kim et al., 2021), and context-specific data augmentation (Zhou et al., 2023a; Li et al., 2021).
2.2 Variational Autoencoders
Variational Autoencoders (VAEs) represent a groundbreaking development in the field of generative models in machine learning (Kingma and Welling, 2013). Fundamentally, a VAE is a probabilistic approach that combines aspects of neural networks with statistical modeling, allowing for efficient encoding and generation of complex data distributions (Doersch, 2016; Kingma and Welling, 2013). Unlike traditional autoencoders that merely compress and decompress data, VAEs introduce a probabilistic twist, enabling them to generate new data samples that are similar to the input data (Kingma and Welling, 2013). This ability to model and sample from complex data distributions has found substantial applications in the realm of smart cities. Recent smart city studies have applied VAEs in multiple urban planning and environmental management applications, with the core capabilities of VAEs summarized as the following:
  • 1.
    Missing Data Imputation: VAEs can facilitate the process of replacing missing data in a dataset with substituted values. The goal is to address gaps and poor quality in data to allow for effective analysis and decision-making. Many previous urban science applications employed the VAE to support data imputation and enrichment in the interest of improving the prediction and forecasting in traffic (Chen et al., 2021; Boquet et al., 2020; Pereira et al., 2020), energy (Shen et al., 2021; Liu and Zhang, 2021), and urban water resources management (Zhang et al., 2023; Wang et al., 2023b; Xie et al., 2019).
  • 2.
    Anomaly Detection: the process involves identifying unusual patterns or outliers in city-related datasets that deviate significantly from the norm. This process is crucial in urban contexts, as it aids in recognizing potential issues or abnormalities in various sectors. VAEs are utilized to improve the Detection of anomalies in complex urban data collected from IoT devices and smart infrastructure (Villegas-Ch et al., 2023; Vu et al., 2020; Mavikumbure et al., 2022).
The versatility and effectiveness of VAEs in data enrichment make them an invaluable tool in advancing smart city initiatives and improving the availability and quality for developing data-driven solutions for sustainable and efficient urban living.
2.3 Transformer-Based Models
Transformer-based models, a groundbreaking innovation in natural language processing (NLP), have revolutionized how machines understand and generate human language (Gillioz et al., 2020). Characterized by their unique architecture, which relies on self-attention mechanisms, transformers facilitate the processing of data sequences, such as text, in a manner that captures contextual relationships more effectively than prior methods (Vaswani et al., 2017). This architecture has been instrumental in the development of advanced language models like ChatGPT, which demonstrates exceptional capabilities in generating human-like text, answering questions, and even assisting in coding and analytics tasks (Feng et al., 2023b; Roumeliotis and Tselikas, 2023).
In recent years, the scientific research community has shown a growing interest in transformer-based models, especially the chatGPT, primarily due to their versatility and efficiency in understanding and contextualizing text for data and knowledge discovery, as well as their intelligent capability for generating codes in multiple programming languages (Sohail et al., 2023). In the urban science sector, GPT and its Application Programming Interface (API) have been used to support a variety of innovative applications in the following aspects:
  • 1.
    Data Mining: analyzing and extracting text-data and image-data from social-media and crowed-sourcing platforms (Hu et al., 2023; Kheiri and Karimi, 2023).
  • 2.
    Autonomous Expert System: leveraging the transformer-based models’ capability to automate domain-specific data discovery (Li and Zhang, 2023) and code generation (Li and Ning, 2023; Feng et al., 2023b), and facilitate decision-support (Singh, 2023; Tan et al., 2023) for complex environmental and urban problems.
2.4 Generative Diffusion Models
Diffusion Models, a class of generative models that have gained prominence in recent years, represent a novel approach in the field of machine learning for synthesizing high-quality data (Ho et al., 2020; Sohl-Dickstein et al., 2015). These models operate on the principle of gradually transforming a distribution of random noise into a distribution of structured data by mimicking the process of diffusion in physics and have broken the long-time dominance of GANs (Yang et al., 2023). The roots of Diffusion Models can be traced back to earlier works in denoising and score-based generative models. Still, the research in the late 2010s truly harnessed their potential in generating complex data such as images and audio. In scientific research, diffusion Models hold significant promise due to their ability to generate realistic and varied data samples, improve likelihood estimation, and handle data with special structures (Yang et al., 2023; Cao et al., 2022). In the urban science and management sector, the recently emerging generative diffusion models are increasingly applied in smart applications to facilitate the simulation and modeling of urban dynamics and processes (Zhou et al., 2023b; Zhong et al., 2023), as well as generative urban design and optimization (Du et al., 2023; Su et al., 2023; He et al., 2023).
Generative diffusion models are also effective in data augmentation (Reutov, 2023) for generating smoothly varied versions of the original data and supporting anomaly detection and denoising. However, their applications in these areas are still rare in the urban science sector. The main strength of diffusion models is their ability to generate intricate, high-fidelity representations of complex systems, making them an invaluable tool for smart city development’s predictive, autonomous planning and optimization.
3 Generative AI in Smart City Applications
This section will provide a structured review of how the recent advancements in generative AI are revolutionizing key areas of urban science research and smart city applications, including traffic and transportation management, mobility research, building operations, energy management, and infrastructure planning. We primarily focus on the applications of the Generative AI models, including GANs, VAEs, transformer-based models, and generative diffusion models. The following subsections will delve into specific application areas of Generative AI models in each of the mentioned smart city areas (as depicted in Figure 2) to highlight their transformative ability to address challenges in urban digital twin applications.
Figure 2: Generative AI application areas in major smart city sectors.
3.1 Data Augmentation
Data augmentation refers to techniques used to increase the amount of data by adding slightly modified copies of already existing data or newly created synthetic data from existing data (Lemley et al., 2017). This is particularly common in machine learning and deep learning, where more training data leads to better model performance. Data augmentation is primarily about creating additional data points to improve the robustness and accuracy of models in tasks like image recognition, classification, and predictions (Moreno-Barea et al., 2020). Generative Adversarial Networks (GANs) and Variational Autoencoders (VAEs) are prevalent in smart city applications for data augmentation, given their proficiency in generating high-quality synthetic data. However, generative diffusion models, known for their strengths in denoising and data augmentation, are primarily utilized in medical research rather than urban science. This suggests a potential untapped area for these models in urban applications, where their unique capabilities could offer novel solutions to complex urban challenges.
3.1.1 Mobility and Transportation Management
Generative AI plays a crucial role in transforming transportation and urban mobility management by leveraging data augmentation techniques. It enhances the accuracy and reliability of predictive models, enabling more efficient traffic flow, route optimization, and smart city planning by generating realistic, varied datasets that mimic complex urban environments.
Zhou et al. (2021) introduces the Trajectory Generative Adversarial Network (TGAN) to address challenges in identifying and linking human mobility patterns, particularly in location-based social networks (LBSNs). By combining adversarial learning and spatio-temporal data analysis, TGAN significantly enhances trajectory-user linking (TUL) by generating synthetic trajectories that closely mimic real human movement patterns. This approach effectively overcomes issues of data sparsity and imbalanced label distribution in trajectory data, demonstrating superior accuracy in TUL compared to existing methods. The use of GANs for data augmentation in this context marks a substantial advancement in human mobility analysis, offering promising implications for various applications, including urban planning and personalized services. Jilani et al. (2022) presents an innovative approach to traffic congestion analysis using GANs for data augmentation. The research introduces a five-layer convolutional neural network model that leverages GAN-enhanced datasets for improved accuracy in classifying traffic congestion. This method significantly improves identifying congestion patterns, achieving a 98.63% accuracy rate. The study highlights the efficacy of GANs in generating realistic synthetic images, thereby enriching the dataset for more effective traffic management and planning applications. Dabboussi (2023) explores an innovative approach using GANs for data augmentation in traffic management. The study presents a novel GAN-based model that successfully generates synthetic traffic data, addressing the critical issue of data scarcity in Intelligent Transportation Systems (ITS). This method demonstrates a significant improvement in traffic prediction models, enhancing their performance through the addition of realistic, GAN-generated data. The research contributes to developing more intelligent and sustainable transportation systems, underlining the potential of GANs in transforming ITS applications. Islam and Abdel-Aty (2021) presents a novel data augmentation technique using a VAE for improving transportation mode classification with limited training data. The VAE’s decoder is adept at generating synthetic data that closely resembles real smartphone sensor data, effectively enlarging the dataset for machine learning models. This method enhances classification accuracy and addresses the imbalance in datasets by boosting underrepresented classes. Its application in activity recognition using smartphone sensors holds significant potential for vehicle-to-pedestrian (V2P) systems and smart mobility solutions. Islam (2021) presents novel deep learning applications in urban transportation systems. It showcases the use of VAEs and GANs for augmenting data in scenarios with limited datasets. The study highlights the capability of these deep learning models to generate synthetic data, which closely resembles real-world data, thus enhancing model training and predictive performance in transportation applications like crash data analysis and traffic pattern prediction. Islam et al. (2021) introduces a VAE-based data augmentation technique to balance crash and non-crash event datasets in traffic systems. Due to the extreme imbalance in these datasets, traditional machine learning models tend to underperform. The VAE encodes events into a latent space, enabling the generation of synthetic crash data that statistically mirrors real data. This method outperforms traditional oversampling techniques like SMOTE and ADASYN, improving specificity and sensitivity in crash prediction models. It also addresses overfitting issues, showcasing VAE’s capability in generating precise, balanced datasets for effective machine learning applications in traffic safety.
3.1.2 Urban Energy Systems
Data augmentation techniques, aided by generative AI, are increasingly pivotal in building energy modeling and power grid applications. These innovative methods have significantly enhanced the accuracy of predictions and forecasting, thereby streamlining the management of urban energy systems.
Wu et al. (2021) presents a novel approach to forecasting energy consumption using a GAN-enhanced ensemble model. This method combines GANs with ensemble learning, demonstrating significant improvements in forecasting accuracy. The study explores the use of GANs for data augmentation, which effectively enriches the training dataset with high-quality, diverse samples. This approach enhances the model’s robustness and generalization capabilities, showcasing the utility of GANs in improving predictive models in energy management systems. The Boicea et al. (2022) explores using GANs for augmenting data in power load forecasting. It demonstrates GAN’s capability to expand a dataset, enhancing the predictive accuracy of a Convolutional Neural Network model used for load forecasting. This approach addresses the challenge of limited sample size in power load forecasting, showing GANs’ effectiveness in data augmentation and their potential in improving forecasting accuracy in energy management systems. Fan et al. (2022) utilizes deep learning, specifically deep generative modeling, for data augmentation in building energy prediction. This approach significantly improves the accuracy of short-term energy predictions in buildings. The study leverages the capabilities of deep learning to generate synthetic but realistic energy data, addressing the challenge of limited high-quality training data. This innovation demonstrates the potential of deep learning in enhancing data-driven models for smart building operation management, especially in scenarios with data scarcity.
3.1.3 Building and Infrastructure Management
Data augmentation, facilitated by generative AI, significantly enhances the predictive analytics and decision-making processes in building and infrastructure management (Baduge et al., 2022). Its application spans a broad spectrum, from structural detection to maintenance forecasting.
Abdollahi et al. (2020) presents a groundbreaking approach in urban planning by integrating GANs, Genetic Optimization Algorithms (GOAs), and Geographic Information Systems (GIS). This study underscores the capability of deep learning, particularly GANs, in generating complex, high-dimensional urban building images. The innovative use of GANs for data augmentation, combined with the optimization strength of GOAs and the spatial analysis of GIS, demonstrates a novel framework in urban planning, enhancing efficiency, scalability, and decision-making in urban development. Li and Betti (2023) addresses a significant challenge in the Structural Health Monitoring (SHM) of civil infrastructure using machine learning (ML), where a lack of comprehensive datasets for training models under various structural damage conditions is a common issue. To overcome this, the research introduces a novel data augmentation strategy within a vibration-based SHM framework, employing a Conditional Variational Autoencoder (CVAE). The CVAE is used to generate additional data samples of power cepstral coefficients of structural acceleration responses in diverse damage scenarios. This is achieved by training the CVAE to model different statistical distributions of these coefficients, using newly defined conditional variables. Notably, the decoder of the CVAE, once trained, is then used independently for the generation of new samples, enhancing the data augmentation process. Hoeiness et al. (2021) utilizes augmented GANs for predicting wind flow in urban environments. This innovative approach rephrases the problem of 3D flow fields into a 2D image-to-image translation challenge. The study explores the use of GANs, particularly Pix2Pix and CycleGAN, along with U-Net autoencoders, demonstrating GAN’s ability to model complex spatial relationships and generate accurate wind flow predictions. This research represents a significant advancement in urban planning and pedestrian comfort analysis for improving street and infrastructure design.
By generating comprehensive datasets that replicate a vast array of environmental and usage conditions, these AI models empower the development of robust solutions, including innovative designs and optimization strategies. Their applications lead to more resilient and efficient operations and maintenance of buildings and urban infrastructure.
3.2 Data synthesis and Scenario Generation
Synthetic data generation involves artificially creating entirely new data that mimics real-world data and is different from data augmentation, which modifies existing data to create variations for enhancing the size and diversity of data sets for training data models (Figueira and Vaz, 2022). Popular generative AI models like GANs and VAEs are instrumental in synthetic data generation for smart city and urban research, offering diverse, realistic datasets where real data is scarce or sensitive due to privacy issues. In many smart city applications, synthetic data supports scenario generation and analysis, enabling planners to simulate and predict outcomes of future and hypothetical urban developments, traffic patterns, and public services in a controlled and risk-free environment, thereby informing more effective and sustainable urban strategies. This approach leads to a more informed and effective scenario-based planning approach, facilitating the evaluation of potential impacts and outcomes of urban policies or changes before they are implemented in the real world.
3.2.1 Mobility and Transportation Management
In human mobility studies, generative models are widely used to generate synthetic trajectories that can realistically reproduce mobility patterns, as well as to enable predictions of future mobility patterns. Luca et al. (2021) have provided a comprehensive survey on the generative AI applications for creating synthetic trajectories. As examples, Smolyak et al. (2020) innovatively uses GANs for anomaly detection in human mobility data. The authors introduce the Infinite Gaussian Mixture Model (IGMM) coupled with bidirectional GANs (BiGANs), an approach that generates realistic synthetic human mobility data and effectively detects multimodal anomalies. This method marks a significant advancement in anomaly detection, addressing the challenges of multimodal data patterns in human mobility and improving the accuracy of identifying anomalous behaviors or trajectories. (Kosaraju et al., 2019) presents a novel approach to pedestrian trajectory forecasting. It integrates a graph-based GAN with a graph attention network (GAT) and Bicycle-GAN to generate realistic multimodal trajectory predictions as future scenarios. This method effectively models the social interactions of pedestrians and accounts for the multimodal nature of human movement. The research demonstrates significant advancements in trajectory forecasting, contributing to applications like autonomous vehicle navigation and urban planning. (Huang et al., 2019) utilizes a Sequential Variational Autoencoder (SVAE) for reconstructing human mobility trajectories. This approach is novel in its integration of a Variational Autoencoder (VAE) with a sequence-to-sequence model, enabling the efficient capture of salient features in human mobility data. The SVAE model demonstrates the capability of VAEs in approximating complex, high-dimensional distributions like urban mobility patterns. This innovation is crucial for enhancing trajectory reconstruction in urban planning and intelligent transportation systems, addressing challenges of data sparsity and privacy concerns. Bao et al. (2020) introduces an innovative approach using Conditional GANs to estimate human mobility during the COVID-19 pandemic. This research significantly contributes by addressing the challenge of estimating human mobility in rapidly changing social contexts with limited training data. The study employs a domain-constraint correction layer within the GAN’s generator, enhancing the learning process and improving the quality of generated mobility predictions. This method demonstrates the potential of GANs in generating reliable spatiotemporal mobility data under uncertain and dynamically changing conditions, a crucial advancement in predictive modeling during public health crises. Wang et al. (2020) introduces SeqST-GAN, a model utilizing Sequence-to-Sequence (Seq2Seq) Generative Adversarial Networks for predicting urban crowd flow. It uniquely addresses the complex spatial-temporal correlations in urban environments by treating crowd flow data as sequences, akin to frames in a video. The combination of Seq2Seq and GANs enables the model to generate accurate multi-step predictions, overcoming the challenge of blurry predictions often seen in deep learning models. This innovative use of GANs demonstrates their capability in generating sequential, high-dimensional data, marking a significant advancement in urban planning and traffic management.
In many previous efforts that aim to develop intelligent transportation systems for efficient traffic planning and management, generative models are developed to create synthetic data for generating planning and simulation scenarios. As examples, (Wu et al., 2020) explores the use of a Long Short-Term Memory (LSTM) network combined with GAN to generate traffic flow scenarios. This approach leverages the LSTM’s capability in capturing temporal dependencies and GAN’s strength in generating realistic traffic patterns. The innovation lies in accurately modeling and simulating complex traffic dynamics, aiding in traffic management and urban planning. This research highlights the potential of deep learning in enhancing predictive models for urban traffic systems. Demetriou et al. (2023) utilizes GANs to create realistic driving scenarios for autonomous vehicle testing. This research demonstrates the innovative use of GANs in generating complex and diverse driving conditions, aiding in the development and validation of autonomous driving systems. The study highlights GAN’s capability to simulate various road environments and traffic situations, which is crucial for ensuring the safety and reliability of autonomous vehicles in real-world conditions. Mladenov et al. (2022) explores a novel application of GANs in traffic simulation. The study introduces an adversarial variational inference technique for calibrating Origin-Destination (OD) matrices in urban traffic simulators. This method is significant for its ability to mitigate the underdetermination problem in traffic simulation, leveraging GANs to match higher-order moments of field data. This innovative approach enhances the accuracy and robustness of traffic simulations, which is essential for effective urban traffic management and planning.
3.2.2 Urban Analytics
Generative AI models excel in creating realistic, high-dimensional data and data-driven scenarios that can mimic complex urban dynamics and processes based on realistic and historic data. By creating synthetic data and scenarios of various urban growth patterns, demographic distributions, and city layouts, these AI tools provide invaluable insights for urban planning, policy making, and morphology analysis, enabling researchers to explore diverse scenarios and predict future urban transformations effectively (De Sabbata et al., 2023).
In terms of detail applications, Johnsen et al. (2022) explores the use of GANs for synthesizing realistic population datasets. This approach leverages GAN’s capability to generate high-dimensional, diverse data, addressing the challenge of limited real-world demographic data availability. By creating synthetic populations that closely resemble actual demographics, the study showcases GANs’ potential in urban planning and policy making, where accurate population models are crucial. Kempinska and Murcio (2019) explores the use of deep learning, specifically VAEs, in modeling urban street networks. It demonstrates how VAEs can effectively capture key urban network metrics in low-dimensional vectors and generate complex urban forms. This approach, capitalizing on the deep learning capability of generating realistic images, marks significant progress in data-driven urban modeling, leveraging the increasing availability of street network data to enhance the quantitative analysis of urban patterns. Zhang et al. (2022) introduces Metropolitan GAN (MetroGAN), a novel Generative Adversarial Network framework, for simulating urban morphology with geographical knowledge. MetroGAN addresses limitations of previous GAN models such as data sparsity and training instability by incorporating a progressive growing structure and a geographical loss for water area constraints. It significantly outperforms existing urban simulation methods, demonstrating over 20% improvement in all metrics. Interestingly, MetroGAN can generate city shapes using solely physical geography features, showcasing its ability to handle various urban attributes and stabilize GAN-based urban simulation. Allen-Dumas et al. (2022) delves into the application of deep learning, particularly GANs, for analyzing urban morphology. This research highlights the capability of GANs to process and interpret complex urban data, providing insightful analyses of city structures and patterns. The innovative use of GANs in this context offers significant contributions to urban planning and development, enhancing our understanding of urban spaces and their evolution. Sun et al. (2021) explores land-use and land-cover change (LUCC) prediction using deep learning. It introduces a model employing GANs, specifically the pix2pix GAN, combined with an attention mechanism. This approach significantly enhances local spatial detail and diversity in urban planning and LUCC prediction. The model’s unique capability lies in fusing multiscale local spatial information with city planning data to predict future land-use changes, showcasing a novel application of GANs in urban planning and environmental monitoring. Kastner and Dogan (2023) presents an innovative GAN model for predicting urban wind flow. This GAN, designed for image-to-image translation, efficiently maps input urban geometry to predicted wind flow patterns. By utilizing Pix2Pix GAN, integrated with a voxelization process and distance-to-object encoding, the model offers rapid and accurate predictions of urban wind flow. This approach significantly reduces the computational cost and time associated with traditional Computational Fluid Dynamics (CFD) simulations, making it a promising tool for urban design and microclimate analysis.
3.2.3 Urban Energy System
Synthetic data supports the planning, management, and operation of urban energy systems by providing realistic, diverse datasets for modeling and simulating various energy scenarios(Li et al., 2020; Roth et al., 2020). These datasets can facilitate the optimization of energy distribution, demand prediction, and efficient resource allocation, reducing carbon-emission in cities and improving efficiency in grid systems.
As detailed examples, Wang and Hong (2020) introduces an innovative approach to generate realistic building electrical load profiles. Utilizing GAN, the research addresses the challenge of generating load profiles that accurately reflect dynamic and stochastic behaviors of real buildings. The study showcases GAN’s capability to anonymize smart meter data, maintaining important statistical information while ensuring privacy. This advancement offers significant implications for smart meter data analytics, enhancing the integration of buildings with the electric grid and promoting sustainable energy management. Dong et al. (2022) discusses a novel data-driven method for generating renewable energy production scenarios using controllable Generative Adversarial Networks (GAN). It focuses on addressing the challenges in power system planning and operation due to uncertainties in wind and solar energy. The proposed GAN model learns the stochastic and dynamic characteristics of renewable resources, enabling the creation of controllable, interpretable scenarios. This method enhances renewable energy modeling, offering significant contributions to the efficiency and reliability of power systems with high renewable penetration. Pan et al. (2019a) introduces an innovative application of deep learning in the field of electric vehicle (EV) load profiles generation. It employs a Variational Auto-Encoder (VAE), a deep learning model, to create EV charging load profiles. This approach is significant because it moves away from traditional methods that heavily rely on probabilistic models and manual data sampling. The VAE model, by learning the inherent characteristics of EV load data, generates diverse load profiles that accurately reflect temporal correlations and probability distribution of actual loads. This methodology signifies a noteworthy leap in using deep learning for more efficient and accurate energy management in the context of increasing EV usage. Jiang et al. (2018) presents a novel approach for wind power scenario generation using GANs. This research stands out for its improvement of GANs, particularly in enforcing Lipschitz continuity through gradient penalty and consistency terms. These enhancements address common issues like overfitting and training instability, leading to more efficient and stable model training. The study demonstrates the effectiveness of this approach in generating high-quality scenarios for wind power, offering significant potential for renewable energy modeling and decision-making under uncertainty. Chen et al. (2018) introduces an innovative approach for renewable energy scenario generation using GANs. This study highlights the ability of GANs to create realistic, data-driven scenarios, capturing the spatial and temporal correlations of renewable power plants. The method is particularly notable for its efficiency and scalability, offering a significant advancement in renewable energy modeling, especially for wind and solar power generation. This work demonstrates the potential of GANs in effectively handling the complexities and variabilities in renewable energy data.
3.2.4 Urban Disaster Management
Generative models, adept at synthesizing realistic images and scenarios, are being utilized for detecting particular disaster and simulating potential disaster impacts, enhancing preparedness strategies. In earthquake and flood management, for instance, GANs generate detailed images of damaged infrastructure, aiding in rapid damage assessment and efficient resource allocation (Tilon et al., 2020). This technology significantly improves response times and decision-making accuracy, proving vital in mitigating the effects of urban disasters and bolstering resilience in city planning and emergency responses.
As detailed examples, Hofmann and Schüttrumpf (2021) presents a novel deep learning approach for real-time pluvial flood prediction. The study introduces floodGAN, a deep convolutional generative adversarial network, which significantly accelerates flood prediction while maintaining high accuracy. This approach effectively utilizes the capabilities of deep learning, particularly GANs, to understand and predict complex flood dynamics based on spatially distributed rainfall events. The innovation lies in its ability to provide fast, accurate flood predictions essential for real-time applications such as early warning systems, marking a substantial advancement in flood modeling and risk management. Salih et al. (2022) explores the use of GANs and Mask Region-Based Convolutional Neural Networks (Mask-R-CNN) for floodwater detection and segmentation in images. The study highlights the capability of GANs to generate segmented images of flooded areas, emphasizing their potential in applications where quick and accurate floodwater identification is crucial. The research demonstrates the effectiveness of GANs in image-based flood detection, offering a novel approach to enhancing flood response and management strategies. Yadav et al. (2024) introduces an innovative unsupervised approach for flood mapping using Synthetic Aperture Radar (SAR) time series data. It leverages a Variational Autoencoder (VAE) trained with reconstruction and contrastive learning techniques. This model is adept at detecting changes by examining the latent feature distributions between pre-flood and post-flood data. Its performance, evaluated against nine different flood events, demonstrates superior accuracy over existing unsupervised models. This approach marks a significant advancement in the application of deep learning to environmental monitoring, specifically in enhancing flood detection capabilities using satellite data. Lago et al. (2023) introduces cGAN-Flood, a conditional generative adversarial network developed to enhance flood prediction in urban catchments. Unlike traditional ANN models, cGAN-Flood can generalize flood predictions to areas outside its training dataset, addressing varying boundary conditions. It achieves this by using two cGAN models for identifying wet cells and estimating water depths, trained with HEC-RAS outputs. The method was tested on different urban catchments and compared with HEC-RAS and the rapid flood model WCA2D. It successfully predicted water depths, despite some underestimation in channels during low-intensity events. Significantly faster than existing models, cGAN-Flood offers a promising, efficient alternative for large-scale flood forecasting. Delacruz (2020) presents an innovative approach using GANs for rapid classification of earthquake-induced damage to railways and roads. It highlights the use of GANs to generate synthetic images of damage, addressing the challenge of limited real-world data for training deep learning models. This method demonstrates the potential of GANs in enhancing the efficiency and accuracy of structural damage assessment, which is crucial for emergency response and infrastructure recovery efforts. Zhou et al. (2023a) utilizes GANs for urban morphology prediction related to local climate zones (LCZs). It employs GANs to integrate LCZ and urban morphology data from six cities, demonstrating superior accuracy and convergence capabilities, particularly with the Pix2pix model. This approach allows for rapid and responsive 3D urban morphology prediction, offering significant insights for enhancing local climates in urban areas through 3D generative design techniques.
3.2.5 Urban Water Management
Currently, the applications of generative AI in urban water management is relatively limited, especially when compared to its broader use in other urban sectors. The primary focus of existing applications lies in the prediction and structural management of sewer water systems. This section will selectively explore and discuss a handful of pertinent examples in this domain. Koochali et al. explores the innovative usages of GANs in urban water management, specifically in improving combined sewer overflow predictions, such as depth and flow rate. GANs are used for data augmentation, generating synthetic time series to balance datasets, thus enhancing the accuracy of prediction models. The study demonstrates that models trained with synthetic data generated by GANs can perform comparably to those trained with real data. The study proposed novel deep-learning approaches that can address challenges like data scarcity in urban water management. Fu et al. (2022) focuses on the application of GANs for tasks like predicting water system behaviors, improving operational efficiency, and enhancing decision-making processes. The paper underlines the effectiveness of GANs in handling complex, data-rich scenarios typical in urban water systems, demonstrating how these advanced AI models can contribute to more sustainable and effective water management strategies. Situ et al. (2021) presents a novel approach for sewer defects detection. It utilizes state-of-the-art Style-based Generative Adversarial Networks (StyleGANs) to generate high-quality synthetic images of sewer defects. These images are then used to train and fine-tune seven well-known CNN models for automated sewer defects detection. The study demonstrates that both variants of StyleGANs are efficient in producing diverse and detailed images, significantly enhancing the performance of deep learning models in sewer defects detection.
3.3 3D City Modeling
The rapid advancement of City Digital Twins is propelled by digital infrastructure and ICT technologies linked to the Internet of Things (IoT), leading to the transformation of traditional cities into smart cities with enhanced infrastructure maintenance, socio-economic growth, and urban governance (Wang et al., 2021; Caprari et al., 2022). Among the majority of the existing digital urban twins, 3D city models has played a pivotal role to support the 3D modelling, monitoring and visualization of the smart city (Masoumi et al., 2023). 3D city models are also developed to enable the interactive and immersive visualization of complex urban environments to facilitate public engagement and collaborative urban planning (Dembski et al., 2019; Xu et al., 2023a), simulation and evaulation of smart technologies (Xu et al., 2023c, d), and 3D cadastral and urban design (Shojaei et al., 2013; Polys et al., 2018).
Despite that the traditional geomatics and procedure-based methods (Lerma et al., 2010; Bejleri et al., 2021) are capable of creating the Digital Surface Models (DSMs) of many large cities (Singh et al., 2013), these DSMs have limited resolution (e.g., texture and geometry) for representing building and infrastructure at the acceptable Level of Details (LoD). Additionally, many novel uses of 3D digital models (e.g., Hardware-In-the-Loop (HIL) simulations and scientific gamification) would require a gamified environment in which urban entities, such as buildings, roads, and infrastructure, are fully separated and have individual physical and gaming properties (Xu et al., 2023c, a). In this setting, creating a high-fidelity 3D representation of a city is a challenging effort that often requires massive, repetitive, time-consuming and manual efforts for surveying and 3D modeling of urban entities (Xu et al., 2023a), along with advanced computing resources for processing and integrating large-scale GIS, remote-sensing (e.g., photogrammetry and LIDAR scan), and crowd-sourcing data in a virtual environment (Xu et al., 2023c). It has been an emerging trend to leverage the 3D content-generation capability of the recent generative AI techniques to automate the tremendous manual efforts for creating 3D assets of urban entities, such as terrain, buildings, infrastructures, and vegetation (Xu et al., 2023c).
3.4 3D Building Generation
The autonomous capability of recent generative AI models has allowed them to create photo-realistic 3D representations of buildings and street environments(Pang and Biljecki, 2022; Wu et al., 2022). These applications have advanced the technical capability of many urban science fields, including architecture design, building construction and operations, and urban planning. Du et al. (2020b) introduces a novel method for 3D building fabrication using a Hybrid GAN. This method incorporates a unique loss function, including cycle consistency and perceptual loss, and utilizes a multi-properties GAN chain to create complex architectural designs. It also features a mixed GAN network for generating geometry and texture coordination. This approach significantly enhances the realism of 3D building models, achieving a 20% performance improvement over traditional methods, demonstrating its potential in applications like digital city construction and smart urban planning. (Kelly et al., 2018) discusses an innovative framework called FrankenGAN. This system is designed to enhance coarse building mass models by adding detailed geometric and texture details, creating realistic and visually appealing urban landscapes. FrankenGAN employs a cascade of Generative Adversarial Networks (GANs), synchronized to maintain consistent styles across multiple scales in neighborhoods. It offers user control over the variability of output styles through exemplar images, allowing for a blend of automated and user-guided design. This technology is significant for its potential to revolutionize urban modeling and architectural design, providing a powerful tool for researchers and professionals in these fields to generate detailed, large-scale 3D city models efficiently and with high fidelity. Bacharidis et al. (2020) explores the use of GANs for reconstructing 3D building facades from images. It employs deep learning techniques for facade structural element detection, depth point-cloud generation, and protrusion estimation. The study demonstrates GANs’ capability in segmenting building components from images and generating detailed 3D facades. These capabilities show GANs’ potential in evaluating urban scenes and extracting urban design parameters, such as window ratio, skyline, and other applications, including cultural heritage preservation. The approach aims to address the challenges of architectural diversity and data scarcity in 3D modeling.
3.5 3D Street Environment Construction
A few other studies have developed generative AI strategy to digitize the entire street environment (with buildings and infrastructure) as 3D models, laying an foundation for the next-generation extended reality applications for supporting participatory and collaborator urban design and planning (Xu et al., 2023a). Tancik et al. (2022) presents a significant advancement in neural rendering for large-scale urban environments. It introduces Block-NeRF, a variant of Neural Radiance Fields (NeRF), specifically designed to handle city-scale scenes. This method effectively decomposes a large scene into individually trained NeRFs, allowing for scalable rendering and enabling updates to specific parts of the environment without retraining the entire network. This technique represents a major step forward in the field of photorealistic 3D reconstruction and neural rendering. (Kim et al., 2020b) introduces a novel method for generating three-dimensional virtual models of cities from a single street-view image. This approach differs from traditional reconstruction methods by utilizing machine learning, specifically generative adversarial networks (GANs) and convolutional neural networks (CNNs), to create terrain maps and identify city components and styles. The generated 3D virtual cities are visually plausible and resemble actual cities. This method represents a significant advancement in the field of urban modeling, offering a unique and efficient way to create detailed, large-scale city models, which is particularly useful for applications in gaming, film, and urban planning. The use of GANs and CNNs in this context is notable for its ability to support research and development in these areas by providing a more streamlined and user-friendly approach to city modeling.
3.6 Generative Urban Design and Optimization
Generative design is defined as using computer-aided generative methods to produce design options, practices, and solutions, improving certain stages of the typical design process (Stojanovski et al., 2020; BuHamdan et al., 2021). In recent years, generative AI models play a crucial role as the intelligent agent in computer-aided urban design. They are employed in many studies of urban design alternatives and strategies that are driven by urban big data and previous solutions (Al Qeisi and Al-Alwan, 2021). The trend of publications that apply generative AI in urban design and optimization has been summarized in Figure 3. Recent progress in generative urban design is evident in the urban science and smart city sector, where these models are increasingly used for tasks such as urban layout and land use optimization, environmental impact assessments, and urban resource management. At a high level, a study by Quan et al. (2019) highlighted the effectiveness of AI in managing urban complexity and sustainability, as well as supporting generative and participatory urban design. Furthermore, the integration of AI in smart cities is demonstrated by the work of Batty (2018), who emphasized AI’s role in analyzing large-scale urban data to inform better city planning and management. The trend towards AI-aided urban planning signifies a paradigm shift, promising smarter, more efficient, and sustainable urban environments.
Figure 3: Summary of generative urban design and optimization publications using Scopus database (accessed on Jan 10th 2024)
As examples of detailed generative urban design applications, Huang et al. (2022) explores the use of GANs, in urban design for environmental performance optimization. It demonstrates the use of GANs in rapidly generating urban design alternatives that meet specific environmental criteria, highlighting deep learning’s ability to process complex spatial data and simulate various urban development scenarios. This approach represents a significant innovation in integrating environmental considerations into the urban design process using advanced AI techniques. Fedorova (2021) examines the use of deep learning, particularly GANs, in creating urban layout designs. It focuses on GANs’ ability to generate diverse and realistic urban layouts, which can significantly aid in urban planning and development. This research underscores the potential of deep learning in addressing complex spatial design challenges, providing innovative tools for urban designers and planners. Noyman and Larson (2020) explores the application of generative AI models, such as GANs and VAEs, in urban design, focusing on generating urban block layouts. This innovative approach does not require predefined parameters for design solutions, as the neural network learns the visual features of existing city structures. The research demonstrates GANs’ ability to adapt designs to various urban contexts and morphologies. Significantly, it highlights the potential of generative AI in urban planning, offering a novel method for designing urban blocks that structurally align with the existing urban fabric without the need for explicit parameter definitions. It should be mentioned that social, economic, and environmental considerations can be used as constraints to improve the urban blocks from the livability and sustainability perspectives. Quan (2022) develops Urban-GAN, an AI-aided generative urban design computation system, which empowers public participation in urban design through Generative Adversarial Networks (GANs) and Deep Convolutional Neural Networks (DCNNs). This innovative system enables individuals with little design expertise to generate professional-level urban designs. It employs sub-symbolic representation and case-based reasoning, facilitating users to select urban form cases and generate similar design schemes. The system’s application in hypothetical design experiments across cities like Manhattan, Portland, and Shanghai illustrates its effectiveness in generating diverse urban forms, highlighting its potential in democratizing urban design and planning. KIM et al. (2020) focuses on developing an AI tool called PlacemakingAI, which uses Generative Adversarial Networks (GANs) for urban design. The tool facilitates public participation in urban planning, allowing users to visualize sustainable urban spaces in real-time. It leverages both supervised and unsupervised GANs to generate synthetic images from datasets of walkable streets. This research is significant for its innovative use of GANs in urban design, enabling a participatory approach and fostering engagement from citizens, designers, and stakeholders in the urban design process.
4 Future of Generative AI in Urban Digital Twins
This section explores the integration of generative AI models with emerging urban digital twin applications (e.g., web-based platforms, cyber-physical systems, and cyberinfrastructure), aiming to augment their current capabilities for more efficient smart city management. Additionally, we delve into the potential technical and methodological challenges associated with the development of generative urban digital twins, underscoring the complexities and innovations required for this visionary approach.
4.1 Cognitive Digital Twins: Opportunities
As an integration between AI and digital twins, cognitive digital twins represent a significant advancement in the digital twin concept, integrating enhanced capabilities beyond their traditional counterparts. These advanced models assimilate cognitive functions, enabling them to dynamically process and learn from data. They harness deep learning to analyze complex datasets, predict future scenarios, and make decisions autonomously, leading to improved efficiency and adaptability. In urban science and smart city applications, cognitive digital twins are proving invaluable for enabling the real-time situational awareness and predictive analytics of complex urban systems using a variety of neuron networks for classification, faulty detection, object detection, and optimization purposes. They also facilitate human-AI collaboration by involving users (e.g., experts, policymakers, and city residents) in the urban management processes through visual dashboards and visualizations. Nevertheless, cognitive digital twin applications that leverage the capability of generative AI to solve complex urban problems are currently limited in the smart city sector.
Figure 4: Potential opportunities of generative AI techniques for enhancing cognitive urban digital twins at the conceptual level.
The vision of incorporating generative AI techniques in cognitive urban digital twins is to achieve an autonomous and generative system that will be cognitive, context-aware, fully automated, and cost-effective. This paper believes that the capability of previous generative AI applications reviewed in the previous section can be seamlessly integrated into cognitive urban digital twins. Future directions or opportunities for the generative urban digital twins are discussed as follows (as depicted in Figure 4).
Enhancing Situational Awareness:
Through data augmentation, generative AI techniques can generate realistic urban data to refine the spatiotemporal resolutions and facilitate anomaly detection in the raw data, such as point clouds and satellite images (Wu et al., 2022). Therefore, these techniques enhance the urban digital twins’ capability to replicate the dynamics and reflect the state of the real-world urban systems in the digital environment.
Improving Prediction Accuracy:
With improved data quantity, quality, and realistic synthetic data that are generated based on real-world data, the urban digital twins’ predictive analytics capability can be improved to predict or forecast the dynamics and response of different urban systems to either futuristic or hypothetical scenarios through data and physics-driven models. For instance, the design and development of critical infrastructure such as airports and their complex interactions with urban transportation systems can be modeled and simulated through different scenarios by integrating proposed future developments (e.g., buildings and transport networks) with current and real-time data (e.g.traffic, current buildings, and land uses). These ”what-if” scenarios also provide alternatives to digital twin users for more adaptive decision support.
Facilitating Human-AI Collaboration:
By generating visualizations, text, and data-driven insights that are understandable to humans, generative AI models can facilitate better collaboration between humans and AI systems. These models can be designed to learn more interpretable features, which are more aligned with human-understandable concepts derived from the domain knowledge. The explanations derived from these models can be made more accessible and sensible to digital twin users.
Enabling Autonomous System Development:
The ability of generative AI models to generate realistic data and codes in different programming languages can be applied to automate the software development of urban digital twins’ components, such as 3D city models and scenario-based simulations (e.g., traffic, mobility, and building energy). This generative ability can significantly increase the adaptability and scalability of digital twin platforms by extending their coverage to new geographic areas and urban domains. For instance, the process of building development approvals, which involves structural, energy, and site examinations, can be automated.
Supporting Decision Optimization:
Generative AI models can generate counterfactuals based on multiple ”what-if” scenarios. These generated urban design or planning strategies can be used as alternatives to help urban planners evaluate the feasibility and efficiency of numerous planning or design alternatives and select the optimized solutions. This addresses the current gap in comprehensive impact analysis of urban design scenarios. This is due to the limitation of computational models and algorithms. For instance, in a Transit Oriented Development (TOD) scenario, different urban design analytics, including building height, sky view factor, walkability, and accessibility to different modes of transport, need to be evaluated.
Promoting AI-aided Design:
Generative models can rapidly generate a wide range of urban design concepts based on specific criteria or policy (e.g., walkability, accessibility, safety, and security, climate resilience), helping planners explore creative ideas and inspirations that might not be immediately visible to humans. Additionally, these models can tailor and modify existing human designs to meet a community’s or environment’s specific needs, and suggest design elements that are more likely to be successful in a particular area.
4.2 Generative AI-Digital Twin Integration
Figure 5: Technical strategies for AI-Digital Twin integration: Generative (Gen)-AI models deployment in different digital twin architecture.
Through the evolution history of the urban digital twins, a few software architectures emerge as particularly adept at fostering sophisticated, scalable, and efficient digital representations of urban landscapes. Among these, monolithic architecture, microservices, and edge-computing architecture are pivotal (Ferko et al., 2022). Monolithic architecture lays the groundwork by facilitating interoperable services that communicate across a network and is crucial for developing basic cyberinfrastructure for amalgamating diverse simulation systems and data sources using the traditional server-client paradigm (Gos and Zabierowski, 2020). Evolving from monolithic architecture, microservices architecture advances this model by organizing applications into collections of loosely coupled services that are often adopted in large-scale cloud-based applications. This structure promotes agile development and deployment of complex cyberinfrastructure systems, enabling seamless updates and scalability for specific elements (e.g., simulation models for specific urban subsystems) of an urban digital twin without impacting the overall framework (Ponce et al., 2019; Xu et al., 2023b). Meanwhile, Edge computing introduces a move towards decentralized and distributed data processing, allocating computational tasks nearer to the data origins at the network’s edge nodes (Yu et al., 2017). Examples of these edge nodes are often edge servers that are directly connected to sensors and mobile IoT devices for sensing various types of urban data (Xu et al., 2022; Marjanović et al., 2018). This shift is vital for urban digital twin development, as it allows for real-time data analysis and immediate response critical for the effective management of dynamic urban activities. Collectively, these architectural strategies create a formidable foundation for constructing urban digital twins, delivering the adaptability, scalability, and immediacy required to simulate and manage complex urban systems accurately.
To effectively integrate generative AI-powered applications within urban digital twins—often manifested through web-based platforms, cyberinfrastructures, and cloud computing services for enhanced smart city functionalities—we propose several technical strategies (as demonstrated in Figure 5). These strategies are designed to bolster scalability, performance, and precision in the AI-digital twin synergy. Tailored specifically to accommodate urban digital twins, these approaches are aligned with distinct architectural patterns. This ensures that each strategy not only addresses the unique challenges of integrating AI into these complex systems but also leverages the inherent strengths of the chosen architectural framework. By doing so, our approach facilitates the creation of dynamic, responsive urban digital twins that can evolve with the changing needs of smart city ecosystems.
Generative AI as a Service:
To integrate AI models into digital twins that adopt monolithic architecture and microservices architecture, a reliable practice would be hosting these models as web services following the paradigm of the service-oriented architecture (Schnicke et al., 2020). Currently, many cloud computing platforms provide the foundational infrastructure needed for deploying ML and DL models as web services at scale, offering flexible resources and cyber-deliver capability that can dynamically adjust to the computational demands of generative AI. Utilizing the recent containerization technologies, such as Docker, and orchestration tools like Kubernetes can further enhance scalability and manageability by allowing AI components to be deployed as microservices. This modular approach facilitates the seamless integration of generative AI into existing urban digital twin frameworks, ensuring that complex data-driven simulations and data analyses run efficiently across distributed computing environments.
Ad-hoc Data Processing and Analytics at Edge Nodes:
For urban digital twin applications that involve real-time data streaming and simulation capabilities, deploying generative AI models on the edge nodes can reduce latency and bandwidth use by processing data closer to the data source (e.g., sensors and IoT devices), which is crucial for real-time applications like traffic flow optimization, building and grid operations, or emergency response simulations within urban digital twins. Generative AI models can be readily deployed on edge servers or data loggers that are directly connected to urban sensors for processing and distributing raw data. This practice is complemented by the use of data streaming and real-time analytics platforms, such as Apache Kafka and Apache Spark, which can handle the high-velocity data feeds sensed from urban environments and provide the necessary processing power to augment and enrich urban data for supporting enhanced situational awareness and simulation-based predictive analytics in real-time.
GPT-powered Autonomous System:
The integration of generative pre-trained transformer (GPT) models into software tools is becoming increasingly popular, supporting research across various scientific domains. Recent studies have demonstrated the potential of leveraging these models to develop autonomous systems and AI agents (Li and Ning, 2023; Vemprala et al., 2023). Such advancements aid in research and decision-making processes, addressing complex scientific challenges by combining GPT APIs with domain-specific data and simulation models. By employing Python or JavaScript APIs provided by GPT, researchers can seamlessly integrate generative AI capabilities into urban digital twins’ various components. This integration facilitates data integration, augmentation, and enrichment, enhancing AI-assisted urban design, education, decision-making, and the interpretation of intricate urban data and simulation models. This approach not only streamlines the incorporation of advanced AI technologies into urban planning and management but also opens new avenues for innovative, data-driven solutions to urban challenges.
4.3 Challenges and Limitations
Although numerous generative AI approaches have been developed to address challenges across different sectors of smart cities, significant obstacles persist. These challenges are often linked to the limitations of the generative AI models themselves or stem from technical issues related to integrating AI with urban digital twins.
Inherent Model Instability:
Generative AI models, such as GANs, often face difficulties in the data training processes to make these models reach a stable state where the generator and discriminator are well-balanced. This instability can manifest in several ways, including mode collapse, non-convergence, vanishing gradients, hyper-parameter sensitivity, and divergence (Sajeeda and Hossain, 2022; Saxena and Cao, 2021).
Evaluation of AI-Generated Data:
The quality of AI-generated data is influenced by several complex factors, including the absence of objective evaluation criteria, challenges in detecting mode collapse, the lack of benchmark data, and issues arising from imbalanced data (Saxena and Cao, 2021). Furthermore, there is a notable lack of consensus regarding assessment indicators for various types of AI-generated data, such as time series and spatial data. This ambiguity in evaluation standards can be attributed partly to the limited number of published studies in this area, which restricts the development of universally accepted benchmarks and metrics.
Computing, Data Storage, and Latency:
In the context of smart cities and urban digital twin applications, generative AI introduces substantial computational challenges, necessitating high processing power to facilitate complex simulations of urban systems and real-time data analytics. Moreover, these applications demand rapid and efficient data storage solutions to handle the extensive data volumes generated and processed, encompassing everything from sensor data to detailed environmental models. The seamless integration of these technologies requires robust, scalable computing infrastructures and sophisticated data storage systems, designed to support the relentless and intensive data streams inherent to smart urban environments.
Network Latency:
The network latency associated with the data exchange for sending the data to generative AI models must be minimized to support real-time smart city functionalities that perform time-sensitive analytics and cyber-physical controls, such as traffic management optimization and incident detection. Addressing these demands often entails equipping urban digital twins with ample communication bandwidth and computational resources to process these AI models in real time, ensuring efficient and effective smart city operations. In fact, generative AI models, such as the VAE, can also be used for data compression, potentially addressing the network latency issues.
5 Conclusion
This survey investigates and summarizes the recent progress on popular generative AI models and their applications in the smart city domain. To enhance the reliability, performance, and efficiency of urban planning and management, the augmentation and synthesis of urban data based on previously collected and simulated data receive a lot of attention in the transportation, urban mobility, energy systems, infrastructure, and urban design sectors. Generative AI models support these sectors’ urban research and management efforts through their excel capabilities for data augmentation, synthetic data and scenario generation, 3D city modeling, and generative urban design. With the advent of urban digital twin technologies, we envision an integration between the generative AI models and cognitive digital twins for a more reliable, autonomous, scalable, and cost-effective management and digital transformation of complex urban systems. Based on a structured review of the previous generative AI application in the smart city sector, this survey outlines 6 promising opportunities for the proposed generative AI-enabled urban digital twins and outlines 3 technical strategies to facilitate the generative AI-digital twin integration within different software engineering architectures. Finally, the survey closes with a discussion of 4 potential challenges that may present in the AI-digital twin integration processes. We believe the integration of generative AI models could significantly revolutionize the existing urban digital twin applications towards more intelligent, reliable, and autonomous systems for supporting smart city development and operations.